# Learning stochastic dynamics from snapshots through regularized unbalanced optimal transport

### v1

@@ 2.1 Schrödinger Bridge Problem
$\mathcal{D}_{\text{KL}}\left(\mu^{\boldsymbol{X}}_{[0,1]}||\mu^{\boldsymbol{Y}}_{[0,1]}\right)=\mathbb{E}_{\mathbb{P}}\left[\mathcal{E}_{1}\left(\int_{0}^{1}\gamma({\boldsymbol{Y}}_{t},t)\mathrm{d}{\boldsymbol{W}}_{t}-\frac{1}{2}\int_{0}^{1}\left\|\gamma({\boldsymbol{Y}_{t}},t)\right\|_{2}^{2}\mathrm{d}t\right)\right],$ (2.6)
@@
Combine with Eq. (2.6), we obtain that
$\mathcal{D}_{\text{KL}}(\mu^{\boldsymbol{X}}_{[0,1]}||\mu^{\boldsymbol{Y}}_{[0,1]})$ $=\mathbb{E}\left[\mathcal{E}_{1}\left(\int_{0}^{1}\frac{1}{2}\left\|\gamma(\boldsymbol{Y}_{t},t)\right\|_{2}^{2}\mathrm{d}t\right)\right]$
$=\mathbb{E}\int_{0}^{1}\frac{1}{2}\left\|\gamma(\boldsymbol{X}_{t},t)\right\|_{2}^{2}\mathrm{d}t$
@@
The proof is done.
We refer to Eq. (2.4) and Eq. (2.5) as the dynamic diffusion Schrödinger bridge.
2.2 Optimal Transport
The optimal transport problem has been extensively studied in various fields. Given two probability distributions $\alpha\in\mathbb{R}_{+}^{n}$ and $\beta\in\mathbb{R}_{+}^{m}$, its primary goal is to find the optimal coupling $\pi\in\mathbb{R}_{+}^{n\times m}$ to transport a given distribution of mass (or resources) from one location to another while minimizing the cost associated with the transportation. The static optimal transport problem can be mathematically formulated as:
$\min_{\pi\in\boldsymbol{\Pi}(\alpha,\beta)}\langle\pi,c\rangle,\quad\boldsymbol{\Pi}(\alpha,\beta)=\left\{\pi\in\mathbb{R}_{+}^{n\times m}:\pi\mathbf{1}_{m}=\alpha,\pi^{T}\mathbf{1}_{n}=\beta,\pi\geq 0\right\}.$
The cost matrix $c\in\mathbb{R}^{n\times m}$ defines the transportation cost between each pair of points, where $c_{ij}:=c(x_{i},y_{j})$ represents the cost of moving a unit mass from point $x_{i}$ to point $y_{j}$. We refer to (Peyré et al., 2019) for more details. Next we briefly stated some well-known results on the dynamical formulation of the OT.
Dynamical Optimal Transport
The formulation is also known as the Benamou-Brenier formulation (Benamou and Brenier, 2000), which can be stated as follows.
Compared to SB, in OT the distributions are connected by a deterministic transport equation instead of diffusion. It can be shown that this dynamical formulation corresponds to a static Kantorovich’s optimal transport problem with cost function $c(x,y)=\|x-y\|_{2}^{2}$.
Regularized Optimal Transport
The regularized optimal transport is defined by a dynamical form of the general Schrödinger bridge by taking the diffusion rate as constant and scaled by $\sigma^{2}$ in Eq. (2.4)).
It can be demonstrated that as $\sigma^{2}$ approaches zero, the solution to this problem converges to that of the Benamou-Brenier problem (Mikami and Thieullen, 2008).
Unbalanced Optimal Transport
To account for unnormalized marginal distributions and effects such as growth and death, an unbalanced optimal transport problem with Wasserstein–Fisher–Rao (WFR) metric has been proposed (Chizat et al., 2018a, b) or its extensions (Gangbo et al., 2019). Here we adopt the WFR unbalanced optimal transport.
[…]
Here $g(x,t)$ is a scalar function that denotes the growth or death rate of particles at the state $x$ and time $t$, and is also optimized in the total energy term. $\alpha$ is the hyperparameter of weight. We should also note that in this case, $\nu_{0}$ and $\nu_{1}$ are not necessarily the normalized probability densities, but are generally densities of masses.
3 Regularized Unbalanced Optimal Transport
In this section, we first introduce the RUOT problem, then derive the Fisher regularization form of RUOT, and discuss its connections with existing literature.
3.1 RUOT with Fisher Regularization
To simplify the notation and illustrate a commonly used setting, we take a special case considering $\boldsymbol{a}(\boldsymbol{x},t)=\sigma^{2}(t)\boldsymbol{I}$, and the general case is left to Section 3.2 for further discussion. Inspired by unbalanced dynamic optimal transport, the dynamical formulation Theorem 2.1 suggests a natural approach to relaxing the mass conservation constraint by introducing a growth/death term $gp$ in Eq. (2.5) as
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\sigma^{2}(t)\boldsymbol{I}p\right)+gp.$
Meanwhile, we also define a loss functional in Eq. (3.7) which incorporates the growth penalization and Wasserstein metric considered in Eq. (2.4) (with a rescaling by $\sigma^{2}(t)$). We refer to this formulation as the regularized unbalanced optimal transport.
Definition 3.1 (Regularized unbalanced optimal transport).
Consider
$\inf_{\left(p,\boldsymbol{b},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\boldsymbol{b}(\boldsymbol{x},t)\right\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t+\int_{0}^{1}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\boldsymbol{x},t)\right){p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (3.7)
where the infimum is taken all pairs $\left(p,\boldsymbol{b}\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,T)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\sigma^{2}(t)\boldsymbol{I}p\right)+gp.$ (3.8)
Note that in the definition if $g(\boldsymbol{x},t)=0$, $\Psi(g)=+\infty$ unless $g=0$ and $\Psi(0)=0$ it degenerates to the regularized optimal transport problem. If $\sigma(t)\rightarrow 0$ and $\Psi\left(g(\boldsymbol{x},t)\right)=|g(\boldsymbol{x},t)|^{2}$ this degenerates to the unbalanced dynamic optimal transport with Wasserstein-Fisher-Rao (WFR) metrics. Meanwhile when $\sigma(t)$ is constant, it coincides with the definition of RUOT provided in (Baradat and Lavenant, 2021). Then, we can reformulate Definition 3.1 with the following Fisher information regularization.
Theorem 3.2.
The regularized unbalanced optimal transport problem (3.7) is equivalent to
$\inf_{\left(p,\boldsymbol{v},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}-\frac{\sigma^{2}(t)}{2}\left(1+\log p\right)g+\alpha\Psi\left(g\right)\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (3.9)
where the infimum is taken all pairs $\left(p,\boldsymbol{v},g\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,T)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)+g(\boldsymbol{x},t)p.$ (3.10)
Proof 3.3.
From (3.8) we obtain
@@ 3.1 RUOT with Fisher Regularization
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}+\frac{1}{2}\left\langle\boldsymbol{v}(\boldsymbol{x},t),\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p\right\rangle+\alpha\Psi\left(g\right)\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t.$
Letting $H\left(p\right):=\int p\log p$ be the entropy, we have
$H\left(p_{1}\right)-H\left(p_{0}\right)$ $=\int_{0}^{1}\partial_{t}H\left({p}(\boldsymbol{x},t)\right)\mathrm{d}t$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\left(1+\log{p}(\boldsymbol{x},t)\right)\partial_{t}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t$
@@
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\left\langle\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p,\boldsymbol{v}(\boldsymbol{x},t)\right\rangle\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t$ $=H\left(p_{1}\right)-H\left(p_{0}\right)$
@@
Remark 3.4.
When $g=0$ and $\Psi(0)=0$, then Eq. (3.9) is same with the dynamic entropy-regularized optimal transport form as discussed in (Bunne et al., 2023a; Li et al., 2020; Gentil et al., 2017; Pooladian and Niles-Weed, 2024; Neklyudov et al., 2024; Léger and Li, 2021)
Remark 3.5.
An additional term in Eq. (3.9)
$\mathcal{I}(p)=\int_{\mathbb{R}^{d}}\|\nabla_{\boldsymbol{x}}\log{p}(\boldsymbol{x},t)\|_{2}^{2}p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}$
is referred to as the Fisher information. Notably, when considering growth/death factors, Eq. (3.9) includes not only the Fisher-Rao metric but also an additional cross-term $\int_{\mathbb{R}^{d}}-\frac{1}{2}\sigma^{2}(t)\left(1+\log{p}(\boldsymbol{x},t)\right)g(\boldsymbol{x},t)p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}.$
Remark 3.6.
In (Buze and Duong, 2023), the authors defined a RUOT problem with nonlinear Fokker-Planck equation constraints. In Appendix A, we show that the proposed form is indeed consistent with the RUOT defined here, when $\sigma(t)$ is constant and $\Psi\left({g}(\boldsymbol{x},t)\right)$ take the quadratic form i.e. $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$ It has been conjectured that the proposed form is the higher order approximation scheme to the unbalanced optimal transport problem as $\sigma\rightarrow 0$.
From the Fokker-Plank equation and the proof of Theorem 3.2, the original SDE $\mathrm{d}{\boldsymbol{X}}_{t}=\left(\boldsymbol{b}\left({\boldsymbol{X}}_{t},t\right)\right)\mathrm{d}t+{\sigma}\left(t\right)\boldsymbol{I}\mathrm{d}{\boldsymbol{W}}_{t}$ can be transformed into the probability flow ODE
Conversely, if the probability flow ODE’s drift $\boldsymbol{v}(\boldsymbol{x},t)$ the diffusion rate $\sigma(t)$ and the score function $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t)$ are known, then the the SDE’s drift term $\boldsymbol{b}(\boldsymbol{x},t)$ can be determined by
$\boldsymbol{b}(\boldsymbol{x},t)=\boldsymbol{v}(\boldsymbol{x},t)+\frac{1}{2}\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t).$
So to specify an SDE is equivalent to specify the probability flow ODE and its score function.
3.2 RUOT in the General Case
For the general case, we consider the following equation
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t)p\right)+gp,$ (3.11)
with the corresponding metric tensor
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\boldsymbol{b}(\boldsymbol{x},t)\right\|_{2}^{2}p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t+\int_{0}^{1}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\boldsymbol{x},t)\right)p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (3.12)
Definition 3.7 (Regularized unbalanced optimal transport II).
Consider
$\inf_{\left(p,\boldsymbol{b},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\boldsymbol{b}(\boldsymbol{x},t)\right\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t+\int_{0}^{1}\int_{\mathbb{R}^{d}}\alpha\Psi\left(g(\boldsymbol{x},t)\right){p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (3.13)
where the infimum is taken all pairs $\left(p,\boldsymbol{b}\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,T)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{b}\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t)p\right)+gp.$ (3.14)
We similarly can reformulate Definition 3.7 with the following Fisher information regularization.
Theorem 3.8.
@@ 3.2 RUOT in the General Case
Consider the regularized unbalanced optimal transport problem Definition 3.7, it is equivalent to
$\inf_{\left(p,\boldsymbol{v},g\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t+\frac{1}{8}\left\|\boldsymbol{\sigma}(\boldsymbol{x},t)\nabla_{\boldsymbol{x}}\cdot\log\left(\boldsymbol{a}(\boldsymbol{x},t)p\right)\right\|_{2}^{2}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t$ (3.15)
where the infimum is taken all pairs $\left(p,\boldsymbol{v},g\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,T)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and
$\partial_{t}p=-\nabla_{\boldsymbol{x}}\cdot\left(p\boldsymbol{v}(\boldsymbol{x},t)\right)+g(\boldsymbol{x},t)p.$ (3.16)
The proof is left to Appendix B. From the proof of Theorem 3.8, the original SDE $\mathrm{d}{\boldsymbol{X}}_{t}=\left(\boldsymbol{b}\left({\boldsymbol{X}}_{t},t\right)\right)\mathrm{d}t+\boldsymbol{\sigma}\left(\boldsymbol{X}_{t},t\right)\mathrm{d}{\boldsymbol{W}}_{t}$ can be transformed into the probability flow ODE
So to specify an SDE is equivalent to specify the probability flow ODE and its score function.
4 Connections with Schrödinger Bridge
Although the RUOT problem can be derived from the dynamical formulation of a Schrödinger bridge relaxation, the relationship between the RUOT problem and the original static Schrödinger bridge problem, the choice of the growth penalty form in RUOT as well as its microscopic interpretation, remains unsolved. In this section, we will discuss two potential microscopic interpretations with the choice of the growth penalty form and identify the problems and challenges associated with each.
4.1 Connections with Branching Schrödinger Bridge
In (Baradat and Lavenant, 2021), the authors explore the relationship between the RUOT and the branching Schrödinger bridge, with the corresponding reference process identified as branching Brownian motion. We will begin by reviewing their principal findings and subsequently highlight the associated issues and challenges.
A Branching Brownian motion (BBM) can be characterized by its diffusivity $\varepsilon$, branching rate $\lambda$ (where each particle is associated with an independent exponential distribution with parameter $\lambda$ ), and the law of offspring denoted by $p=(p_{k})_{k\in\mathbb{N}}\in\mathcal{P}(\mathbb{N})$ (The probability measures on $\mathbb{N}$ with nonnegative sequences summing to one), where $p_{k}$ is the probability of producing $k$ descendants at a branching event. We define $q_{k}=\lambda p_{k}$ and introduce the generating function of $\boldsymbol{q}$ as $\Phi_{\boldsymbol{q}}$.
$\Phi_{\boldsymbol{q}}:z\mapsto\sum_{k\in\mathbb{N}}q_{k}z^{k}=\lambda\sum_{k\in\mathbb{N}}p_{k}z^{k}.$
The branching mechanism is entirely determined by $\boldsymbol{q}$ , since $\lambda$ and $\boldsymbol{p}$ can be recovered from $\boldsymbol{q}$ by $\lambda=\sum_{k\in\mathbb{N}}q_{k}$ and $p_{k}=\lambda^{-1}q_{k}$ for $k\in\mathbb{N}$ Therefore, these parameters are sufficient to fully characterize the BBM with diffusivity $\nu$ , branching mechanism $\boldsymbol{q}=\lambda\boldsymbol{p}$ , and initial law $R\sim\text{BBM}(\varepsilon,\boldsymbol{q},\nu_{0})$ Their core result indicates that consider Definition 3.1 where $\sigma(t)=\sigma$ is constant and $\varepsilon=\sigma^{2}$ then if the growth penalty term is defined as follows, it can correspond to a branching Schrödinger process.
$\Psi_{\varepsilon,\boldsymbol{q}}^{*}(s)=\varepsilon\left(\Phi_{\boldsymbol{q}}\left(e^{s/\varepsilon}\right)e^{-s/\varepsilon}-\Phi_{\boldsymbol{q}}(1)\right)=\varepsilon\sum_{k=0}^{+\infty}q_{k}\left\{\exp\left((k-1)\frac{s}{\varepsilon}\right)-1\right\},$ (4.17)
and we define $\Psi_{\varepsilon,\boldsymbol{q}}(r)=\sup_{s\in\mathbb{R}}rs-\Psi_{\varepsilon,\boldsymbol{q}}^{*}(s)$. So by following their result, we can present some examples.
The case of only loss mass
This example has been included in their work. We consider $\varepsilon=\lambda=1$, $p_{0}=1$ and $\sum_{k\geq 2}p_{k}=0$, which means the particle can only die at the branching event. Substituting it into Eq. (4.17), we have
$\Psi_{1,\boldsymbol{q}}^{*}(s)=\exp\left(-s\right)-1.$
Then by computing its Legendre transform we obtain
$\Psi_{1,\boldsymbol{q}}(r)=1+r-r\log(-r),$
where $r<0$. This form is consistent with the formula discussed in (Chen et al., 2022b).
The case of only gain mass
$\Psi_{1,\boldsymbol{q}}(r)=1-r+r\log(r),$
which $r>0$.
The case of both gain mass and loss mass
We consider $\varepsilon=\lambda=1$, $p_{0}=p_{2}=\frac{1}{2}$, which means the particle can divide into two parts or die at the same probability at the branching event. This is also presented in their work, which we have
$\Psi_{1,\boldsymbol{q}}^{*}(s)=\cosh(s)-1.$
General case
Although we have illustrated several examples, challenges persist, particularly in the context of more general scenarios. Specifically, when the growth penalty term cannot be expressed as $\Psi_{\varepsilon,\boldsymbol{q}}$, the connection to the corresponding stochastic process remains unclear. For instance, this ambiguity arises in cases where in the RUOT the $\Psi(g)$ takes the quadratic form i.e., the WFR metric, or when $\Psi(g)$ is in a linear form, such as $|g|$. These issues still require specific treatment and to our knowledge, remain an open problem.
4.2 Connections with Schrödinger Bridge on Weighted Path
where $\mu^{\boldsymbol{X}}_{[0,1]}$ represents the probability measure induced by the SDE (Eq. (2.2)). Here for example we consider a simple case when $g(\boldsymbol{x},t)=C$. Then following the procedure of Theorem 2.1.
$\mathcal{D}_{\text{KL}}(\mu^{\boldsymbol{G}}_{[0,1]}||\mu^{\boldsymbol{Y}}_{[0,1]})$ $=\mathbb{E}_{\mu^{\boldsymbol{G}}_{[0,1]}}\left(\int_{0}^{1}\gamma({\boldsymbol{Y}_{t}},t)\mathrm{d}{\boldsymbol{W}}_{t}-\frac{1}{2}\left\|\gamma({\boldsymbol{Y}_{t}},t)\right\|_{2}^{2}\mathrm{d}t+C\right),$ (4.18)
5 Learning RUOT through Deep Learning
Given unnormalized distributions at $T$ discrete time points, $\boldsymbol{X}_{i}\sim\mu_{i}$ for fixed timepoints $i\in\{0,\ldots,T-1\}$, we aim to learn a continuous stochastic dynamic process satisfying the RUOT problem from the data. Similarly, to simplify the exposition and illustrate a commonly used case, we consider Definition 3.1 with the stochastic dynamics
$\mathrm{d}{\boldsymbol{X}}_{t}=\left(\boldsymbol{b}\left({\boldsymbol{X}}_{t},t\right)\right)\mathrm{d}t+{\sigma}\left(t\right)\boldsymbol{I}\mathrm{d}{\boldsymbol{W}}_{t}.$
As previously discussed in Theorem 3.2, we approach this by transforming the problem into learning the drift $\boldsymbol{v}(\boldsymbol{x},t)$ of the probability ODE and its score function $\frac{1}{2}\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x},t)$. We parameterize $\boldsymbol{v}(\boldsymbol{x},t)$, ${g}(\boldsymbol{x},t)$, and $\frac{1}{2}\sigma^{2}(t)\log{p}(\boldsymbol{x},t)$ using neural networks $\boldsymbol{v}_{\theta},~{}g_{\theta}$ and $s_{\theta}$ respectively.
Figure 1: Overview of DeepRUOT.
5.1 Energy Function in Theorem 3.2
To compute equation Eq. (3.9), direct calculation is infeasible due to the high dimensionality. Thus, we need to transform it into an equivalent form that can be evaluated using Monte Carlo methods. Adapting from the approach in (Sha et al., 2024), we have the following theorem.
Theorem 5.1.
If smooth density $p(\boldsymbol{x},t):\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{+}$, velocity field $\boldsymbol{v}(\boldsymbol{x},t):\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ and growth rate $g(x,t):\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}$ satisfy
for all $0\leq t\leq T$ with $\frac{\mathrm{d}\boldsymbol{x}(t)}{\mathrm{d}t}=\boldsymbol{v}(\boldsymbol{x},t)$ and $\boldsymbol{x}(0)=\boldsymbol{x}_{0}$, then for any measurable function $f(\boldsymbol{x},t):\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$, we have
$\int_{0}^{1}\int_{\mathbb{R}^{d}}f(\boldsymbol{x},t)p(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t=\mathbb{E}_{\boldsymbol{x_{0}}\sim p_{0}}\int_{0}^{1}f(\boldsymbol{x},t)e^{\int_{0}^{t}g(\boldsymbol{x},s)\mathrm{d}s}\mathrm{d}t.$
According to this theorem, Eq. (3.9) is equivalent to the following form
@@ 5 Learning RUOT through Deep Learning
$W_{0,T}=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}-\frac{\sigma^{2}(t)}{2}\left(1+\log p\right)g+\alpha\Psi\left(g\right)\right]p\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (5.19)

### v2

@@ D.1 Proof of Dynamical Formulation of Schrödinger Bridge
$\mathcal{D}_{\text{KL}}\left(\mu^{\boldsymbol{X}}_{[0,1]}||\mu^{\boldsymbol{Y}}_{[0,1]}\right)=\mathbb{E}_{\mathbb{P}}\left[\mathcal{E}_{1}\left(\int_{0}^{1}\gamma({\boldsymbol{Y}}_{t},t)\mathrm{d}{\boldsymbol{W}}_{t}-\frac{1}{2}\int_{0}^{1}\left\|\gamma({\boldsymbol{Y}_{t}},t)\right\|_{2}^{2}\mathrm{d}t\right)\right],$ (14)
@@
Combine with Eq. (14), we obtain that
$\mathcal{D}_{\text{KL}}(\mu^{\boldsymbol{X}}_{[0,1]}||\mu^{\boldsymbol{Y}}_{[0,1]})$ $=\mathbb{E}\left[\mathcal{E}_{1}\left(\int_{0}^{1}\frac{1}{2}\left\|\gamma(\boldsymbol{Y}_{t},t)\right\|_{2}^{2}\mathrm{d}t\right)\right],$
$=\mathbb{E}\int_{0}^{1}\frac{1}{2}\left\|\gamma(\boldsymbol{X}_{t},t)\right\|_{2}^{2}\mathrm{d}t,$
@@
The proof is done. ∎
D.2 Proof of Fisher Regularization of RUOT
Proof.
From (6) we obtain
@@ D.2 Proof of Fisher Regularization of RUOT
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}+\frac{1}{2}\left\langle\boldsymbol{v}(\boldsymbol{x},t),\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p\right\rangle+\alpha\Psi\left(g\right)\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t.$ (15)
Letting $H\left({p}(\boldsymbol{x},t)\right):=\int_{\mathbb{R}^{d}}{p}(\boldsymbol{x},t)\log{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}$ be the entropy, we have
$H\left(p(\boldsymbol{x},1)\right)-H\left(p(\boldsymbol{x},0\right)$ $=\int_{0}^{1}\partial_{t}H\left({p}(\boldsymbol{x},t)\right)\mathrm{d}t,$
$=\int_{0}^{1}\int_{\mathbb{R}^{d}}\left(1+\log{p}(\boldsymbol{x},t)\right)\partial_{t}{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$
@@
$\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\left\langle\sigma^{2}(t)\nabla_{\boldsymbol{x}}\log p,\boldsymbol{v}(\boldsymbol{x},t)\right\rangle\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t$ $=\sigma^{2}(t)\left(H\left(p(\boldsymbol{x},1)\right)-H\left(p(\boldsymbol{x},0)\right)\right)$
@@
D.3 RUOT with Nonlinear Fokker-Planck Equation Constraints
Note that when $\Psi\left({g}(\boldsymbol{x},t)\right)$ take the quadratic form i.e. $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$, then by canceling the cross term in Theorem 4.1 we will get another Fisher information regularisation form as below.
Theorem D.1.
@@ D.3 RUOT with Nonlinear Fokker-Planck Equation Constraints
When $\Psi\left({g}(\boldsymbol{x},t)\right)=|{g}(\boldsymbol{x},t)|_{2}^{2}$, the regularized unbalanced optimal transport problem (7) is equivalent to
@@
$\inf_{\left(p,\boldsymbol{v},\widetilde{g}\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\left\|\boldsymbol{v}(\boldsymbol{x},t)\right\|_{2}^{2}+\frac{\sigma^{4}(t)}{8}\left\|\nabla_{\boldsymbol{x}}\log p\right\|_{2}^{2}+\alpha|\widetilde{g}(\boldsymbol{x},t)|^{2}-\frac{\sigma^{4}(t)}{16\alpha}(1+\log p)^{2}\right]p\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (16)
where the infimum is taken all pairs $\left(p,\boldsymbol{v},\widetilde{g}\right)$ such that $p(\cdot,0)=$ $\nu_{0},p(\cdot,1)=\nu_{1},p(\boldsymbol{x},t)$ absolutely continuous, and